# BED-LLM: Intelligent Information Gathering with LLMs and Bayesian Experimental Design

### v1

@@ 4.3 Extracting the belief state
A further advantage of the prior–likelihood construction is that our belief state on $\theta$ can be extracted directly, as it is simply $p_{f}(\theta;h_{t})$. With the data–estimation construction, we would have to estimate a marginal distribution from $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p\left(y_{t+1};[h_{t},x_{t+1}]\right)p\left(\theta;[h_{t},x_{t+1},y_{t+1}]\right)$} by integrating out the synthetic response $p(\theta;h_{t})=\int{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p\left(y_{t+1};[h_{t},x_{t+1}]\right)p\left(\theta;[h_{t},x_{t+1},y_{t+1}]\right)}\;\mathrm{d}y_{t+1}$. Direct access to $p(\theta;h_{t})$ is also important to ensure that our current belief state is independent of the next question $x_{t+1}$, as the marginal obtained from the integral above is unlikely to remain the same for all $x_{t+1}$ in the way we intuitively want it to.
4.4 Alignment between EIG and belief updating procedure
Our ultimate goal is to minimize uncertainty in $\theta$, as measured by its entropy. With this in mind, we can use the expected uncertainty reduction framework of Bickford Smith et al, (2025) to provide insights into how well our EIG formulation and belief updating procedures align.
To simplify discussions, for now we consider the setting where we choose a single question $x$ and obtain a response $y$. Following Bickford Smith et al, (2025), we can think of the “true” optimal design as selecting
Note here that true optimal design has no direct dependency on our current beliefs about $\theta$; it only depends on $p_{\mathrm{true}}(y;x)$ and the hypothetical beliefs we produce for given observed data, $p(\theta;x,y)$. Thus, we can now see that our choice of joint model corresponds to different choices for approximating these quantities. Assuming that the LLM distribution is used directly for the conditional as per Section˜4.2, we thus have that:
The former of these considerations is difficult to control for as we simply do not know the true response distribution and it is hard to say which approach will thus estimate it best (though we can refer to the discussion in Section˜4.1 to determine which best matches our beliefs about the true response distribution). However, we do know upfront how we plan to derive our belief distribution on $\theta$ in practice, so we can use this to guide which joint model we formulate our EIG from. Namely, we see that using
  • The prior–likelihood EIG pairing equates to assuming we will make a Bayesian update to our beliefs on $\theta$ using the likelihood $p_{\mathrm{LLM}}(y;[\theta,x])$;
  • The data–estimation EIG pairing equates to assuming we will make an in-context update to our beliefs on $\theta$, as we are treating $p(\theta;x,y)$ as $p_{\mathrm{LLM}}(\theta;[x,y])$.
For the settings we consider, we expect $p_{f}(\theta;[x,y])$ to generally be better approximated by a Bayesian update than an in–context update, aligning with our decision to use the prior–likelihood formulation. The reasons for this are that a) the filtering often removes a large proportion of the generated samples, especially at later experiment turns, with $p_{\mathrm{LLM}}(\theta;h_{t})$ not fully incorporating information from the history; b) the maintaining of the set of one consistent hypotheses from one turn to the next encourages a more Bayesian behavior, with samples persisting unless contradicted by a new likelihood term; and c) the typical premature over–collapse of $p_{\mathrm{LLM}}(\theta;h_{t})$ to a small number of hypotheses means it is typically unrepresentative of our beliefs.
The analysis here can get a little more complicated when we consider the fact that “one-step optimal” decisions like Equation˜6 may no longer be optimal in the sequential decision–making setting. We provide more discussion on this in Section˜A.3.
@@ 5 Related Work
In this work, we consider the broad problem of sequential information gathering using LLMs—that is, adaptively asking natural-language questions to allow the LLM to rapidly learn about a parameter of interest $\theta$. Several works have explored the baseline ability of LLMs to ask such questions (Zhang et al,, 2024; Li et al, 2025b, )—effectively, the “naive" baseline in our experiments. While these works demonstrate some ability to adaptively construct information-seeking questions, they often fail to extract important information (Li et al, 2025a, ), leading us to more sophisticated approaches.
Ours is also not the first paper to use a targeted model-based approach to information gathering with LLMs: several existing works have previously attempted to choose questions based on estimates of the expected information gain, marginal predictive entropy, or similar quantities (Piriyakulkij et al,, 2023; Hu et al,, 2024; Kobalczyk et al,, 2025; Wang et al,, 2025; Mazzaccara et al,, 2024; Chan et al,, 2025; Cooper et al,, 2025). However, as explained below, these existing methods all make design choices that significantly limit performance. Further, none of them provide comparable insights into how the joint model should be best formulated and updated from the underlying LLM; most of them even fail to directly link their approach and BED itself.
A number of these previous works, including all that also employ a prior–likelihood pairing in their joint model, actually implicitly select queries using predictive entropy in practice (rather than EIG as claimed) because they assume a constant expected likelihood entropy (Kobalczyk et al,, 2025; Mazzaccara et al,, 2024; Cooper et al,, 2025). As we discussed in Section˜3.2.1 and show empirically in Section˜6, this simplification is not generally reasonable—as predictive entropy is not in general a good approximation to EIG—and leads to worse performance. We believe our work to be the first which uses a prior–likelihood model pairing without simplifying the EIG to just the predictive entropy.
All of the other methods have instead relied on models derived from a data–estimation pairing, then either obtaining a sample-based estimate of $\mathbb{E}_{p\left(y_{t+1};[h_{t},x_{t+1}]\right)}\left[\mathrm{H}[p_{\mathrm{LLM}}\left(\theta;[h_{t},x_{t+1},y_{t+1}]\right)]\right]$ (Piriyakulkij et al,, 2023; Hu et al,, 2024; Wang et al,, 2025), or in the case of Chan et al, (2025), an approximation to this entropy based on the size of conformal sets. As discussed in Section˜4, our prior–likelihood model pairing provides significant benefits, particularly in the large or unbounded hypothesis spaces we explore in this paper. To get an approach which is based on the data–estimation model pairing to work, previous works have thus had to restrict the set of allowable hypotheses (i.e. $\theta$) to a small set, typically also restricting the space of allowing outcomes and further making a deterministic likelihood assumption (thereby again implicitly reducing the EIG to predictive entropy). These restrictions significantly limit applicability and performance.
Rather than augmenting a frozen LLM with the ability to estimate utility functions, some works have instead aimed to post-train an LLM to improve its ability to ask questions (Zhang et al,, 2024; Wu et al,, 2025; Andukuri et al,, 2024). Most do not explicitly consider informativeness of questions: Zhang et al, (2024) and Wu et al, (2025) use reinforcement learning techniques to reward generations that quickly lead to the correct answer, and Andukuri et al, (2024) builds on Li et al, 2025b by fine-tuning on successful traces. Mazzaccara et al, (2024) do indirectly incorporate uncertainty: they use predictive entropy to identify informative questions, and then either fine-tune on the highest-entropy question, or perform DPO comparing the highest-entropy question with a lower-entropy question. We do not address fine-tuning in this work, focusing instead on exploring the correct way to formulate BED using LLMs.
As discussed in Section˜4, a key challenge in adapting BED to the LLM setting is in aligning the expected information gain with the actual uncertainties extracted from the LLM after updating. Handa et al, (2024) take a different approach to this problem by using the LLM to generate features for an external conventional Bayesian joint model (in their case, a linear Bradley Terry model), rather than deriving their joint model more directly from the LLM itself. This can be a good choice when we already have a well-specified Bayesian model form for the problem at hand; however, this may be challenging in arbitrarily large and complex hypothesis spaces. In particular, their specific method is not applicable more widely beyond the preference learning context they consider.
@@ 6 Experiments
We now empirically evaluate our proposed BED-LLM method, compare it against relevant baselines, and analyze its properties through targeted ablation studies. We consider two interactive question–answering scenarios that test different types of sequential information acquisition: 1) 20–Questions: a deductive reasoning game requiring identification of a hidden target entity through binary (Yes/No) queries, and 2) Active Preference Elicitation: a task in which the goal is to infer a latent user preference vector via multiple–choice questions.
To evaluate the different questioner methods (i.e., BED-LLM and our baselines), we further need some form of answerer that represents the user or external environment we are trying to learn about. Due to the complications of evaluating with real human participants, we instead conduct our tests with a simulated user based on a separate LLM instance that has been given a ground truth target $\theta^{*}$ (corresponding to the hidden target entity for 20–Questions and a user profile for the active preference elicitation). Note that this answerer LLM does not itself perform any learning—it does not keep previous question–response pairs in context. While we consider both deriving the questioner and answerer from the same LLM and from different LLMs, we emphasize that they are always using separate LLM instances with different prompts and contexts, thus no information is transferred between them except in the questions and responses themselves.
We compare BED-LLM against two baselines:
a) Entropy: selecting questions using the maximum predictive entropy, as done in Kobalczyk et al, (2025), Mazzaccara et al, (2024), and Cooper et al, (2025). For fairness and to ensure a strong baseline, we keep the entropy setup as close to BED-LLM as possible (including prompts, candidate question generation, and our joint model construction with hypothesis filtering), varying only in replacing the EIG estimator with an estimator of the marginal predictive entropy,
$\mathrm{Entropy}_{\theta}(x_{t+1})=-\sum_{y_{t+1}\in\mathcal{Y}}\hat{p}(y_{t+1};[h_{t},x_{t+1}])\log\hat{p}(y_{t+1};[h_{t},x_{t+1}]).$ (7)
We construct an estimator for Equation˜7 analogously to Equation˜5. We also note here that our version of the Entropy baseline produces noticeably stronger performance on the 20–Questions problem than Kobalczyk et al, (2025) report for their own “EIG” approach, demonstrating it efficacy. [^5] Though Kobalczyk et al, (2025) describe their approach as targeting the EIG, the estimator they use (given by their Algorithm 1) actually equates to a sample-based estimator of the marginal predictive entropy instead and is thus comparable with our Entropy baseline. As they do not provide code specifically for the 20–questions experiments or explicit results on their approaches’ success rate, we were unable to provide direct comparisons against their exact setup. However, we believe that it is still clear from the results they do present that their approach is performing worse than the Entropy baseline we do consider. First, the reported ranks they give indicate they are not beating the naive LLM approach with the same regularity as Entropy does. Second, in their rebuttal on OpenReview they explain that they only report ranks and not success rate because “only very few runs would result in a non-zero accuracy”, whereas we see that our Entropy baseline is regularly successful.
b) Naive QA:we prompt the questioner LLM to directly produce an informative next question without explicit hypothesis generation, EIG computation, or filtering. The question is obtained by drawing a single sample at temperature $T=1$ from the questioner LLM. This approach, which was explored in Zhang et al, (2024), relies on the implicit reasoning capabilities of the LLM to select good questions.
6.1 20–Questions
We first evaluate BED-LLM in the classic “20–Questions” game, where the questioner must identify a hidden target $\theta^{*}$ by asking up to 20 Yes/No questions. Here, the hypothesis space of possible $\theta$ is large and not explicitly defined or restricted: it is bounded only by what the LLM can generate. As such, all previous work that has considered using BED-style decision with LLMs either a) cannot be directly applied in our precise problem setup because it requires the questioner to be given access to an explicit, restricted set of possible $\theta$ (Hu et al,, 2024; Piriyakulkij et al,, 2023; Wang et al,, 2025; Chan et al,, 2025), or b) is a variant of our Entropy baseline, due to using marginal predictive entropy in lieu of the EIG (Kobalczyk et al,, 2025; Mazzaccara et al,, 2024; Cooper et al,, 2025).
We evaluate across three distinct datasets each containing a mix of 100 obscure and common targets corresponding to either Animals, Celebrities, or Things. Here, the dataset is just a list of different $\theta^{*}$ that will be individually provided to the answerer to instantiate different problems (e.g. we conduct a trial where $\theta^{*}=$“dog”, then one where $\theta^{*}=$“cat”, etc). The list of targets is never provided to the questioner model to restrict the set of possible hypotheses: the questioner is only prompted that is trying to identify an “animal”, “celebrity”, or “thing”. The datasets are: a) Animals: a set of animal species generated with OpenAI’s o3 model to ensure a diverse mix and balanced taxonomy, see Appendix˜D for details; b) Celebrities: a diverse set of public figures, as used by Zhang et al, (2024); c) Things: a collection of everyday and exotic entities drawn from the web corpus, as used by Zhang et al, (2024); it covers a wide range of categories, from plants and clothing to professions, events, and mythical creatures. We note that the best success rates achieved by Zhang et al, (2024) on Celebrities and Things are substantially below those of both BED-LLM and our Entropy baseline.
Using our sample–then–filter process (c.f. Section˜3.3), we aim to sample at least $N=15$ hypotheses, repeating the cycle up to three times if needed (the exact number of hypotheses can be less than this as it may not be possible to generate sufficient valid hypotheses, especially in later experiment turns). The questioner generates $M=15$ candidate questions to test, $\mathcal{X}^{\text{cand}}$, using the “conditional generation” approach of Section˜3.1 when possible, but falling back on “unconditional generation” if insufficient candidate hypotheses have been generated.
Evaluation
[…]
Table 1: Success rate (%) for 20–Questions at the end of the game. Best result in bold. Success rates are integers because each dataset contains exactly 100 entities. $\pm$ numbers represent the standard error of the mean estimate (this is estimated using $\sqrt{p(1-p)/(n-1)}$ where $p$ is the success percentage and $n$ is the number of datapoints, noting this estimator is positively biased and thus conservative).
Success Rate (%) $\uparrow$
@@
Dataset Model Naive QA Entropy BED-LLM
@@
Figure 2: Success rate against turn for the 20–Questions game. Error bars represent the standard error of the mean estimate across 100 target entities in each dataset. The same model used as both questioner and answerer.
Figure 3: Success rates for the 20–Questions game with different models: the questioner Q and answerer A.
Results
First, we evaluate in the setting where the same LLM is used as both the questioner and the answerer. Table˜1 shows the success rates for the different datasets and a variety of different LLMs. Across all datasets and all LLMs, BED-LLM achieves the highest success rate, significantly outperforming the Entropy and Naive QA baselines. It is particularly notable how the success rate of our principled information-gain strategy is typically more than double that of Naive QA which relies solely on implicit LLM reasoning.
Per-turn success rates are shown in Figure˜2. Over the first few turns, BED-LLM and Entropy perform similarly. However, after around 7-10 turns we consistently see BED-LLM pull away from Entropy. This may simply be because it takes some time for the effect of choosing good questions to take effect (the large space of possible $\theta$ means that no approach can reliably determine the answer in only a few steps), or it may be that the effect of using proper EIG estimates becomes more pronounced as the set of viable hypotheses decreases and it is harder to find useful questions.
Different questioner and answerer
In the experiments above, the same LLM served as both the questioner and the answerer. This setting implicitly assumes that the questioner model has access to the answerer’s response distribution, i.e. that there is no misspecification in the likelihood used by BED-LLM and Entropy or the internal reasoning of Naive QA. However, in practice, the answerer is typically a human whose responses will follow a different distribution to the questioner’s internal model for reasoning about responses. To test the robustness of the various methods to this model misspecification, we therefore also evaluate scenarios where the questioner and answerer are formulated with entirely different LLMs, instead of just different instances of the same LLM. As shown in Figure˜3, the benefit of BED-LLM persists even under this model misspecification; it consistently achieves higher success rates than both baselines across datasets. In contrast, Entropy exhibits a more severe drop in performance in this more realistic setting, often now performing more in line with Naive QA, which itself appears to be quite robust to the misspecification.
Figure 4: Data–estimation ablation using Qwen2.5-72B-Instruct on Celebrities dataset.
Data–estimation ablation
As described in Section˜3, an important design choice is whether we choose the prior–likelihood pairing or the data–estimation pairing for BED-LLM. As an ablation for this critical decision, we also implement a variant of the alternative Data–Estimation method (see Appendix˜C for details). This gives an approach that is similar to those of Hu et al, (2024); Piriyakulkij et al, (2023); Wang et al, (2025); Chan et al, (2025), but which does not require us to assume a single fixed set of possible hypotheses throughout the sequential BED process. We evaluate this alternative approach on the Celebrities dataset using Qwen2.5-72B as both the questioner and answerer. Results are shown in Figure 4. Consistent with our analysis in Section˜4, BED-LLM substantially outperforms the Data–Estimation alternative. Data–Estimation still outperforms the Naive QA baseline, indicating that even this pairing benefits from explicit EIG–driven question selection over purely heuristic question generation, but interestingly does worse than Entropy, emphasizing the importance of estimating our uncertainty in the $y$ space instead of $\theta$ space. These findings reinforce our claim that the choice of joint–model factorization is a critical design decision for sequential BED in the LLM setting.
6.2 Active Preference Elicitation
In our 20–Questions problems, the generated hypotheses were all single, concrete entities (barring any failures of the LLMs) and most reasonable questions had clear–cut answers (e.g. “is Beyoncé a singer?”). In contrast, many problems, especially those related to eliciting information from a user, involve more abstract targets and/or more subjective and less predictable answers. For example, if we are learning about the characteristics or preferences of a user, then it may be difficult to explicitly define a concrete closed set of possible $\theta$, while it is also quite challenging for the LLM to accurately predict user responses and develop appropriate uncertainty estimates.
To study such a scenario, we evaluate BED-LLM on inferring users’ film preferences. Here the target we are trying to gather information about is somewhat abstract, and we have some flexibility in how we define $\theta$ in our joint model. Our chosen setup is to define $\theta$ to be a natural language paragraph describing a user’s taste in films; we refer to this as a persona. To emulate a human user, our answerer is set up to be an LLM prompted to answer questions as if it were a person with a given persona $\theta^{*}$.
To generate a dataset of ground-truth personas, we take a set of $200$ real user ratings from the MovieLens-100K dataset (Harper & Konstan,, 2015), then use an “oracle” LLM (namely, OpenAI’s o3 model) to produce a natural language persona that is consistent with each distinct set of ratings. As was the case for the 20–Questions problems, this dataset is never provided to the questioner and the set of allowed $\theta$ is not constrained.
Because we need the LLM to be able to meaningfully capture uncertainty in the space of responses as discussed in Section˜4.2, we restrict questions to be multiple-choice. Specifically, the questioner is tasked with producing a question along with five possible responses A/B/C/D/E. We then define each $x_{t}$ to be the question coupled with the possible responses, and each $y_{t}$ to be one of the letters A/B/C/D/E to provide a restricted set of tokens over which we can measure entropy. Option E is further constrained to always be “none of the above” so that the answerer is not committed to choosing one of the directly generated choices if none are suitable.
For BED-LLM and Entropy, we compare $M=8$ candidate questions at each turn and we aim to generate at least $N=5$ candidate hypotheses. We use the “unconstrained generation” approach of candidate question generation (c.f. Section˜3.1) as the user profiles can be quite diffuse and we are only generating a small number of possible hypotheses that can be quite easy to split.
We note that data–estimation setup is not at all viable for this problem because the large number of tokens and varying dimensionality of each $\theta$ sample mean that $\mathrm{H}[p_{\mathrm{LLM}}(\theta;[h_{t},x_{t+1},y_{t+1}])]$ is not only infeasible to estimate, but also is not a meaningful measure of uncertainty.
(a) Qwen2.5-72B
(b) GPT-4o-mini
(c) Llama-3.3-70B
(d) GPT-4o
Figure 5: Results for the active preference elicitation with the same LLM used for the questioner and answerer. Shown is the average ratings of 10 recommendations for each of 200 separate simulated users.
Evaluation
To again allow tracking of progress through the experiment, after each turn of the interaction $t$, the questioner generates ten film recommendations, conditioned on $h_{t}$. These recommendations are checked for consistency with prior questions and answers; if any are judged inconsistent then they are removed and additional recommendations are generated. The quality of the film recommendations is then assessed using an “LLM-as-judge” protocol (Zhu et al,, 2025; Trivedi et al,, 2024). Namely, the answerer evaluates each of the 10 films recommended by the questioner, conditioned on the hidden target user profile $\theta^{*}$. It scores each film on a scale of 1 to 5 (in 0.5 increments), based on how well the recommendation aligns with $\theta^{*}$ — this score is output together with a brief justification to increase reliability. We report the mean rating and standard error across 200 users, over 5 question–answer turns.
Results
In Figure˜5, we show the results using the same model for both questioner and answerer. For all models, BED-LLM establishes an advantage over other baselines by turn $t=3$, although the differences in performance between methods are smaller than for the 20–Questions game. On average, Entropy performs similarly to Naive QA.
(a) Q: Qwen2.5-72B. A: GPT-4o-mini.
(b) Q: GPT-4o-mini. A: Qwen2.5-72B.
Figure 6: Results for Active Preference Elicitation with a different models: questioner Q and answerer A.
Different questioner and answerer
Like in the 20–Questions setting, we now also evaluate performance on Active Preference Elicitation in the misspecified model scenario by using a different LLM for the questioner and answerer. Figure˜6 presents the results. While BED-LLM and Entropy seem to be quite robust to the model misspecification, it is now Naive QA that seems to be brittle, exhibiting a drop in performance when using Qwen2.5-72B questioner and GPT-4o-mini answerer.

### v2

@@ 4 On the Specification of $p(\theta,y_{t};h_{t-1},x_{t})$, and its Implications
Extracting the belief state   A further advantage of the prior–likelihood construction is that our belief state on $\theta$ can be extracted directly as $p_{f}(\theta;h_{t-1})$. With the data–estimation construction, we would have to estimate the marginal on $\theta$ by integrating $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p\left(y_{t};[h_{t-1},x_{t}]\right)p\left(\theta;[h_{t-1},x_{t},y_{t}]\right)$} over the synthetic response $y_{t}$. Direct access to $p(\theta;h_{t-1})$ is also important to ensure that our current belief state is independent of the next question $x_{t}$, which is both intuitively desirable and theoretically required to be a valid BED approach (Lindley, 1972); data–estimation formulations will generally violate this.
@@ 5 Related Work
Several works have explored the baseline ability of LLMs to rapidly learn about a parameter of interest by asking questions (Zhang et al., 2024; Li et al., 2025b)—effectively our Naive baseline in §6. While these works demonstrate some ability to adaptively construct information-seeking questions, they often fail to extract important information (Li et al., 2025a).
Some works have further specifically attempted to choose questions based on model-based uncertainty criteria (Piriyakulkij et al., 2023; Hu et al., 2024; Kobalczyk et al., 2025; Mazzaccara et al., 2024; Cooper et al., 2025). None of these works provide the same careful consideration of how the underlying joint model should be formulated, which underpins our own work, and they all assume deterministic likelihood models that mean their objectives correspond to a sample-based estimate of marginal predictive entropy in practice, as explained in Section˜3.2. In general, these previous works have also required restrictions on the space of allowable hypotheses, $\theta$, and typically require additional assumptions and/or approximations. More extensive discussion of related work is given in Appendix˜C.
@@ 6 Experiments
We now assess how well BED-LLM and alternative information-gathering approaches work in two practical scenarios: 20 Questions, a game in which the player has to guess a target entity and can ask up to 20 yes-no questions about the entity; and preference elicitation, a task in which the agent has to predict a user’s preference profile and can ask five multiple-choice questions to the user.
Answerer
We produce answers to the questioner LLM’s questions using a separate answerer LLM. The answerer is provided with a ground-truth $\theta^{*}$ (a target entity in 20 Questions or a user profile in preference elicitation) and processes individual questions from the questioner without access to any of the questioner’s context (i.e. $h_{t-1}$and the questioner’s prompts). We test two questioner-answerer setups, where the two are served by separate instances of the same LLM, or two different LLMs. The latter scenario is important because in practice, the answerer will often follow a different distribution than the questioner’s internal model for reasoning about responses, thereby forming a model misspecification.
Baselines
We compare BED-LLM against two existing baseline methods: Naive and Split. Naive involves prompting the questioner to directly generate an informative next question, without explicit hypotheses generation or computing a data-acquisition objective, and then sampling the question with temperature T=1T\!=\!1; this was explored by Zhang et al. (2024). Split involves choosing the question that most equally splits a sampled set of hypotheses $\Theta^{\text{cand}}$, which corresponds to maximizing the marginal predictive entropy $\mathrm{H}[p(y_{t};x_{t},h_{t-1})]$ in a model with a deterministic likelihood. As such, the methods of Cooper et al. (2025), Hu et al. (2024), Kobalczyk et al. (2025), Mazzaccara et al. (2024) and Piriyakulkij et al. (2023) can all be viewed as variants of this Split baseline. While Split is not applicable to the preference-elicitation scenario, where a deterministic likelihood is not viable, it is to our knowledge the state-of-the-art method for 20 Questions. We note that our own Split baseline implementation achieves dramatically better results than reported by, for example, Kobalczyk et al. (2025), so this constitutes a very strong baseline relative to previous work. On top of this, our implementation of Naive appears to signficantly improve over that of Zhang et al. (2024),
6.1 20 Questions
We consider three sets of 20 Questions problems: Animals, Celebrities, and Things (See Section˜F.1). Each problem set comprises 100 target entities $\{\theta^{*}_{i}\}_{i=1}^{100}$ from a given category. The space of possible $\theta$ is large and not explicitly defined or restricted: we do not tell the LLM this set of target entities, so the space of $\theta$ is bounded only by what the LLM can generate. We note that by comparison, many previous works have relied on restricted spaces for $\theta$ (Chan et al., 2025; Hu et al., 2024; Piriyakulkij et al., 2023; Wang et al., 2025).
Success Rate (%)
@@
Dataset Model Naive Split BED-LLM Entropy Data–Est. ICL Beliefs Impl. Max.
@@
Table 1: Success rate (%) for 20 Questions at the end of the game. Best result in bold. $\pm$ numbers show the standard error of the mean estimated using $\sqrt{p(1-p)/(n-1)}$ where $p$ is the success percentage and $n$ is the number of datapoints. This estimator is positively biased and thus conservative. Data–Est. is not possible to run for GPT models due to limited logprobs support in OpenAI API.
[…]
BED-LLM improves over Naive and Split baselines
Our results in Table˜1 and Figure˜1 show BED-LLM significantly outperforming Naive and Split across all problems and LLMs. Particularly notable is that BED-LLM’s final success rate is typically more than double that of Naive, highlighting the big gains that can be achieved by using explicit EIG maximisation instead of implicit LLM reasoning.
Figure 1: Success rate on 20 Questions: mean $\pm$ standard error across 100 targets per dataset.
BED-LLM ablations
In order to understand the importance of BED-LLM’s algorithmic components, we further evaluate four ablations on this problem: Entropy, Data–Estimation, ICL Beliefs and Implicit Maximization. Each of these differs from BED-LLM with respect to one algorithmic component. Entropy replaces the EIG data-acquisition objective with the marginal predictive entropy $\mathrm{H}[p(y_{t};x_{t},h_{t-1})]$ (Section˜3.2); this contrasts with Split in that it uses $p_{\mathrm{LLM}}(y_{t+1};[\theta,x_{t+1}])$ rather than a deterministic likelihood. Implicit Maximization involves sampling candidate questions and prompting the questioner to select the most informative one, without any explicit objective estimation (Section˜3.2). ICL Beliefs uses $\theta$ belief updates derived from simple in-context learning, namely $p_{\mathrm{LLM}}(\theta;h_{t-1})$ instead of $p_{f}(\theta;h_{t-1})$, testing the importance of our filtering mechanism (Section˜3.3). Data–Estimation uses a data–estimation pairing rather than BED-LLM’s prior–likelihood setup (see Appendix˜D). In Table˜1 we see that BED-LLM comfortably outperforms all alternative approaches. Notably, Entropy provides the strongest ablation baseline, with performance almost identical (sometimes superior) to Split. This shows that the use of a non-deterministic likelihood is beneficial in allowing us to target a proper EIG, rather than because it is particularly detrimental to the marginal predictive entropy itself.
Prior–likelihood outperforms data–estimation
Our analysis in §4 is validated by our results: BED-LLM’s prior-likelihood approach substantially outperforms Data–Estimation. Data–Estimation still outperforms Naive, but interestingly it performs worse than Entropy, highlighting the importance of estimating uncertainty in the $y$ space instead of $\theta$ space. These findings reinforce our claim that the choice of joint-model factorization is a critical algorithmic decision.
Rejection sampling and explicit EIG maximization are key
We also see the importance of two other aspects of BED-LLM. First, how we produce our beliefs over $\theta$ matters: deriving beliefs using simple in-context learning, as in ICL Beliefs, lead to massive performance drops. Second, while BED-LLM’s routines for sampling candidate questions and hypotheses are crucial, they alone are not sufficient: passing the samples to an LLM and prompting it to select the highest-EIG question, as in Implicit Maximization, works much less well than using the samples to explicitly maximize EIG.
BED-LLM is robust to questioner–answerer mismatch
Our results in Figure˜1 demonstrate that the benefit of BED-LLM persists even under model misspecification. This is important for applicability to real-world users, whose responses will follow a different distribution than the questioner LLM.
6.2 Preference Elicitation
Unlike 20 Questions, in which $\theta$ is a concrete entity and most reasonable questions have clear answers, many real-world information-gathering tasks involve more abstract targets and less predictable data generation. A key example is learning user preferences, where it may be difficult to explicitly define a concrete closed set of possible $\theta$, and it is also challenging for the LLM to develop appropriate uncertainty estimates. To study such a scenario, we evaluate BED-LLM on inferring users’ film preferences. Here the target we are trying to gather information about is somewhat abstract, and we have some flexibility in how we define $\theta$ in our joint model. Our chosen setup is to define $\theta$ to be a user profile, namely a paragraph of text describing the user’s film preferences, with our answerer model prompted to emulate a user with a given profile; see Appendix˜G for full details. We consider 200 different user profiles as the ground truth $\theta^{*}$, but as with 20 Questions this set is never given to the questioner. Because Split is not applicable as a baseline here (a deterministic likelihood assumption is clearly unreasonable), we benchmark with the similar Entropy approach instead. We also note that data–estimation setup is completely unviable here as well because of the large $\theta$ space.
At each turn $t\in(0,1,\ldots,5)$ we use the question-answer history as context for generating a list of ten film recommendations. This list is then rated in its fit to the user profile using an LLM-as-judge setup (Trivedi et al., 2024; Zhu et al., 2025). Specifically, the answerer scores each film on a scale of 1 to 5 (in 0.5 increments), based on how well the film aligns with $\theta^{*}$; this score is output together with a brief justification to increase reliability. The films’ scores are not included in $h_{t-1}$.
Our results in Figure˜2 show that, while Naive is often a strong baseline in this preference-elicitation scenario, BED-LLM is still able to provide a boost over both Naive and Entropy, producing higher-rated film recommendations. BED-LLM’s benefit is most clear in scenarios where the questioner belongs to a different model class to the answerer: here Naive’s performance is much less convincing.
Figure 2: Mean rating across 10 film recommendations: mean $\pm$ standard error across 200 users.